# Faster cosmological analysis with power spectrum without simulations

### 31 Jan 2024 v2

@@ 4.2.1 Application to data
Figure 4: (Top:) A heatmap of the logarithmic (base 10) of the compression matrix. Middle: A heatmap of the logarithmic of the absolute value of the compressed covariance matrix. Except for the diagonal at $\sim\log(1)=0$, all the off-diagonal elements are less than any of the elements of the uncompressed matrix. Therefore, the compressed covariance matrix resembles the identity matrix except for some numerical noise. This implies the MOPED compression is successful. Bottom: A heatmap showing the logarithm (base 10) of the absolute difference between the compressed covariance matrix and the identity matrix of the same size. The color bar shows the maximum deviation of the compressed covariance matrix from the identity matrix is on the order of $10^{-11}$. Again, this could be explained by computational numerical error, so the compression is done successfully.
As a first test of the MOPED algorithm, we construct the compressed covariance matrix for our full 6dFGS+BOSS+eBOSS data vector and the analytical covariance matrices from Section 4.1. In total, we have 9 different samples and each sample has 8 different independent free parameters. [^7] We follow d'Amico et al. (2020) and Glanville et al. (2022) to set $c_{r,2}=0$ and let $c_{r,1}$ absorb the contribution of $c_{r,2}$ to the quadrupole when we only fit the monopole and quadrupole. This reduces the number of free bias parameters from 10 to 9. Furthermore, we set $c_{4}=0$. Therefore, we only have 8 free parameters. We also have 4 free cosmological parameters which are shared across all 9 samples. After compression, the size of the model and data power spectrum becomes $8\times 9+4=76$. Compared to its original size of 342 measurement bins, the compressed power spectrum is more than four times smaller.
Lastly, we want to check whether the MOPED compression is successful. From equation (18) and Heavens et al. (2000) we can see that if the compression is successful, the covariance matrix measured from a set of simulations where the data vector of each simulation has had the compression applied should be the same as the identity matrix with a reduced size. Fig 4 illustrates that the absolute difference between the compressed covariance matrix and the identity matrix is everywhere less than $10^{-10}$ and much smaller than the amplitude of the compression matrix itself. Hence we consider the differences between the compressed covariance and identity matrices as due only to the computational numerical error in the calculation and conclude the compression has been done successfully.
@@ 4.3 Using Taylor expansion to interpolate the power spectrum
Recent advancements in machine learning allow us to use emulators to compute the power spectrum much faster than before. In this work, we use one of the simplest forms of emulators, Taylor expansion, to interpolate the likelihood during the data analysis. We adopted the method developed by Colas et al. (2020) for PyBird and they found it does not impact the constraints of cosmological parameters with the BOSS DR12 data set.
It takes three steps to generate the grid of power spectrum for PyBird. Firstly, we want to set the cosmological parameters at the center of the grid. These initial parameters are set according to table 3. We have two different sets of tests: "Template" and "Guess". If we try to use the analytical covariance matrix and MOPED compression for future surveys, we won’t know the best-fit parameters beforehand. However, we can make educated guesses of the best-fit parameters based on previous surveys such as Planck (Planck Collaboration et al., 2020). For the "Template set", the values are taken from the Planck 2018 constraints. For the second set of parameters, they are chosen such that all cosmological parameters except $\Omega_{b}h^{2}$ are more than $3\sigma$ from the mean of the posterior with the simulated covariance matrix. We did not do the same thing for $\Omega_{b}h^{2}$ because the BBN priors for the galaxy surveys typically dominate it. Furthermore, we also need the nuisance parameters to calculate the analytical covariance matrix and the compression matrix. For simplicity, for both "Template" and "Guess", we assume $b_{1}=1.8$ and all other parameters equal to 0.5. During our analysis, we only find $b_{1}$ has a strong influence on the final constraints and it is strongly degenerate with $A_{s}$. This is expected since both change the amplitude of the power spectrum. Furthermore, we can also get $b_{1}$ independently through HOD analysis, so we won’t consider the case where the nuisance parameters are far away from the best-fit. For the second and higher iterations, the input cosmological and nuisance parameters are set to the best-fit of the previous iteration. In this work, the width of each grid cell for cosmological parameters is set to $\Delta\ln{(10^{10}A_{s})}=0.2,\Delta h=0.02,\Delta\Omega_{\mathrm{cdm}}h^{2}=0.01,$ and $\Delta\Omega_{b}h^{2}=0.0005$. For each parameter, we have 9 different grid cells (1 at the center, 4 at the positive/negative direction).
@@
It is worth noting that, from the derivations laid out in Section 4.2, one could alternatively 1) interpolate the un-windowed power spectrum and multiply it by the product of the compression and window function matrices or 2) interpolate the fully compressed model on the grid. We do not implement approach 1) because the derivatives of the windowed power spectrum are required to build the compression matrix, and can be obtained automatically from the Taylor expansion in our adopted approach. Furthermore, the theoretical model vector before convolution is typically evaluated in much narrower bins that the data, such that it would be more expensive to store the grids, and the multiplication of the model with the convolved window would be more computationally expensive. We do not investigate approach 2) in this work because it would require more extensive validation of the Taylor expansion methodology itself (which to our benefit was carried out in Colas et al. 2020), and could be difficult in general as there is no longer the ability to identify a ‘scale’ at which the Taylor expansion may break down. So we leave this to future work.
@@
Table 3: These are the initial cosmological parameters to generate the grid of the power spectrum. There are two different sets of parameters, the parameters from the first set "Template" are not the best-fit parameters but are taken from Planck 2018 (Planck Collaboration et al., 2020) constraints. The parameters in the second set "Guess" are chosen such that the cosmological parameters are more than $3\sigma$ away from the mean of the posterior with the simulated covariance matrix except for $\Omega_{b}h^{2}$. This is because $\Omega_{b}h^{2}$ is dominated by the BBN prior, so it is unlikely it will deviate from its fiducial value.
@@ 5 Result and Discussion
In this section, we will summarize the result from fitting the 6dFGS, BOSS, and eBOSS power spectra with three different covariance matrices (simulation, analytical, and analytical with compression). Here, we will only show the constraints on the cosmological parameters because they are the most important results. In this work, we use the emcee (Foreman-Mackey et al., 2013) package to analyze the data. The convergence criteria for all the fits are set to the default convergence criteria of emcee.
@@
The constraints on the cosmological parameters are shown in Fig. 5 and their respective numerical values in Table 4. Inside the brackets are the best-fit values computed with the basin-hopping algorithm from Scipy (Wales & Doye, 1997). Both Fig. 5 and Table 4 demonstrate the constraints on cosmological parameters using the simulated, analytical, and analytical covariance matrix with compression are in good agreement with each other. For the first column in Table 4, S denotes simulation, C denotes compression, A denotes analytic, 1 denotes the first iteration, and 2 denotes the second iteration. Additionally, G means we used the “Guess" parameters in Table 3 to generate the analytical covariance matrix and the MOPED compression for the first iteration. We treat the constraints with the simulated covariance matrix as the “ground truth".
@@
Going further and also using an analytic covariance, using the “Template" initial parameters in Table 3 to generate the analytical covariance matrix and the MOPED compression, the biases in cosmological parameters are around $0.3\sigma$ ($0.1\sigma$) for the mean of the posterior (best-fit). Furthermore, the cosmological constraints with the analytical covariance matrix from the first iteration are also slightly tighter than the ones with the simulated covariance matrix.
@@
In either case, we then perform a second iteration, generating the analytical covariance matrix and the MOPED compression with the best-fit values from the first iteration. After the second iteraion, table 4 shows the biggest deviation of $0.2\sigma$ arises in the constraint on the Hubble parameter with the analytical covariance matrix. Furthermore, if we also apply the MOPED compression, this is reduced to $0.1\sigma$. This applies regardless of whether we started with the better “Template” parameters or the “Guess” parameters. The mean of posterior for $\Omega_{b}h^{2}$ is generally shifted by around $0.14-0.18\sigma$. However, it is worth noting that the constraint on the baryonic density is quite tight, and dominated by the BBN prior, so we find it generally takes a longer time than other parameters to converge and is also of less interest. We hence believe this to be a sufficient demonstration of the method.
@@ 6 Conclusions
In this work, we combine the analytical covariance matrix, the MOPED compression, and use Taylor expansion to interpolate the power spectrum for the first time to analyze power spectrum data from the 6dFGS, BOSS, and eBOSS surveys. We demonstrate that this removes the necessity to generate large ensembles of simulations when analyzing such data (a potential bottleneck for future surveys such as the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2019) and Euclid (Laureijs et al., 2011)). while also offering potentially significant speed-ups in model fitting.
For the state-of-the-art dataset tested herein and without assuming a priori knowledge of the best-fit parameters, we use two different sets of initial parameters to generate the analytical covariance matrix and the MOPED compression. The “Template" parameters (within $1\sigma$ of the mean of the posterior with the simulated covariance matrix) and the “Guess" parameters (more than $3\sigma$ from the mean of the posterior with the simulated covariance matrix). The analytical covariance matrix takes around a day to compute, which is much faster than a few months for the simulated covariance matrix. Furthermore, the main bottleneck of the covariance matrix calculation is calculating the window kernels and their normalization in Wadekar & Scoccimarro (2020) from the random file. The cosmology-dependent part of the calculation that needs to be recalculated for the second iteration only takes around half an hour. After two iterations, we found the mean of the posteriors and best-fit values of cosmological parameters from these two different sets agree with each other and are consistent with the ones with the simulated covariance matrix within $0.1-0.2\sigma$ regardless of the choice of initial guess. The remaining difference is likely due to the sampling noise.
We noticed the mean of the posterior of cosmological parameters does not necessarily agree with their best-fit values because of the prior volume effect. The best-fit values should be used to generate the analytical covariance matrix and the MOPED compression for the second iteration. Furthermore, the MOPED compression is able to speed up the MCMC by around a factor of four. This is because, with the Taylor expansion, the matrix multiplication becomes the main bottleneck, the MOPED compression can significantly reduce the size of the matrices which speeds up the MCMC.
Further work could investigate whether the compressed data vector itself could be emulated, further decreasing the run time. One could also determine a more rigorous set of convergence criteria for determining how many global iterations to perform, beyond the two we use here, that are suitable for guaranteeing the robustness of results from next generation surveys

### 05 Apr 2024 v3

@@ 4.2.1 Application to data
Figure 4: (Top:) A heatmap of the logarithm (base 10) of the compression matrix. Middle: A heatmap of the logarithm (base 10) of the absolute value of the compressed covariance matrix. Except for the diagonal at $\sim\log(1)=0$, all the off-diagonal elements are close to zero. Therefore, the compressed covariance matrix resembles the identity matrix except for some numerical noise. This implies the MOPED compression is successful. Bottom: A heatmap showing the logarithm (base 10) of the absolute difference between the compressed covariance matrix and the identity matrix of the same size. The color bar shows the maximum deviation of the compressed covariance matrix from the identity matrix is on the order of $10^{-11}$. Again, this could be explained by computational numerical error, so the compression is done successfully.
As a first test of the MOPED algorithm, we construct the compressed covariance matrix for our full 6dFGS+BOSS+eBOSS data vector and the analytical covariance matrices from Section 4.1. In total, we have nine different samples, and each sample has eight different independent free parameters.555see Table 1 We also have four free cosmological parameters shared across all nine samples. After compression, the size of the power spectrum becomes $8\times 9+4=76$. Compared to its original size of 342 measurement bins, the compressed power spectrum is more than four times smaller.
Lastly, we want to check whether the MOPED compression is successful. Equation (21) and Heavens et al. (2000) show that if the compression is successful, the compressed covariance matrix should be the identity matrix. Fig 4 illustrates that the absolute difference between the compressed covariance matrix and the identity matrix is everywhere less than $10^{-11}$ and much smaller than the amplitude of the compression matrix itself. Hence we consider the differences between the compressed covariance and the identity matrix to be due to the numerical computational error. Therefore, the compression is successful.
@@ 4.3 Using Taylor expansion to interpolate the power spectrum
Recent advancements in machine learning allow us to use emulators to compute the power spectrum much faster than the traditional approach. In this work, we use one of the simplest forms of emulators, Taylor expansion, to interpolate the likelihood during the data analysis. This method was developed by Colas et al. (2020) for PyBird. They found it does not impact the constraints of cosmological parameters with the BOSS DR12 data set.
It takes three steps to generate the grid of power spectrum for PyBird. Firstly, we want to set the cosmological parameters at the center of the grid. These initial parameters are set according to table 3. We have two different sets of tests: "Template" and "Guess". If we try to use the analytical covariance matrix and MOPED compression for future surveys, we will not know the best-fit parameters beforehand. However, we can make educated guesses of the best-fit parameters based on previous surveys such as Planck (Planck Collaboration et al., 2020). The values in the “Template" set are taken from the Planck 2018 constraints. For the second set of parameters, they are chosen such that all cosmological parameters except $\Omega_{b}h^{2}$ are more than $3\sigma$ from the mean of the posterior with the simulated covariance matrix. We did not do the same thing for $\Omega_{b}h^{2}$ because the BBN priors for the galaxy surveys typically dominate it. Furthermore, we also need the nuisance parameters to calculate the analytical covariance matrix and the compression matrix. For simplicity, we assume $b_{1}=1.8$ and all other parameters equal to 0.5 for both "Template" and "Guess" sets. During our analysis, we find only $b_{1}$ influences the final constraints because it strongly degenerates with $A_{s}$. This is expected since both change the amplitude of the power spectrum. Furthermore, we can also get $b_{1}$ independently through HOD (Halo Occupation Distribution) analysis. Therefore, we only vary the cosmological parameters, not the nuisance parameters between the two initial conditions. For the second and higher iterations, the input cosmological and nuisance parameters are set to the best-fit of the previous iteration. In this work, the width of each grid cell for cosmological parameters is set to $\Delta\ln{(10^{10}A_{s})}=0.2,\Delta h=0.02,\Delta\Omega_{\mathrm{cdm}}h^{2}=0.01,$ and $\Delta\Omega_{b}h^{2}=0.0005$. For each parameter, we have nine different grid cells (1 at the center, and four in the positive/negative direction).
@@
It is worth noting that, from the derivations laid out in Section 4.2, one could alternatively 1) interpolate the un-windowed power spectrum and multiply it by the product of the compression and window function matrices or 2) interpolate the fully compressed model on the grid. We do not implement approach 1) because the derivatives of the windowed power spectrum are required to build the compression matrix, and can be obtained automatically from the Taylor expansion in our adopted approach. Furthermore, the theoretical model vector before convolution is typically evaluated in much narrower bins than the data, such that it would be more expensive to store the grids, and the multiplication of the model with the convolved window would be more computationally expensive. We do not investigate approach 2) in this work because it would require more extensive validation of the Taylor expansion methodology itself (which to our benefit was carried out in Colas et al. 2020), and could be difficult in general as there is no longer the ability to identify a ‘scale’ at which the Taylor expansion may break down. So we leave this to future work.
@@
Table 3: These are the initial cosmological parameters to generate the grid of the power spectrum. There are two sets of parameters. The parameters from the "Template" set are not the best-fit parameters but are taken from Planck 2018 (Planck Collaboration et al., 2020) constraints. The parameters in the "Guess" set are chosen such that the cosmological parameters are more than $3\sigma$ away from the mean of the posterior with the simulated covariance matrix except for $\Omega_{b}h^{2}$. It is because $\Omega_{b}h^{2}$ is dominated by the BBN prior, so it is unlikely to deviate from its fiducial value.
@@ 5 Result and Discussion
In this section, we will summarize the result from fitting the 6dFGS, BOSS, and eBOSS power spectra with three different covariance matrices (simulation, analytical, and analytical with compression). Here, we will only show the constraints on the cosmological parameters because they are the most important results. In this work, we use the emcee (Foreman-Mackey et al., 2013) package to analyze the data. We use the default affine invariant sampler for the MCMC and follow emcee’s recommendation that the chain is converged when the number of iterations is more than fifty times longer than the integrated auto-correlation time. [^6] See https://emcee.readthedocs.io/en/stable/tutorials/autocorr/ for more detail.
@@
The constraints on the cosmological parameters are shown in Fig. 5 and their respective numerical values in Table 4. Inside the brackets are the best-fit values computed with the basin-hopping algorithm from Scipy (Wales & Doye, 1997). Both Fig. 5 and Table 4 demonstrate the constraints on cosmological parameters using the simulated, analytical, and analytical covariance matrix with compression are in good agreement with one another. For the first column in Table 4, S denotes simulation, C denotes compression, A denotes analytic, 1 denotes the first iteration, and 2 denotes the second iteration. Additionally, G means we used the “Guess" parameters in Table 3 to generate the analytical covariance matrix and the MOPED compression for the first iteration. We treat the constraints with the simulated covariance matrix as the “ground truth".
@@
Going further and also using an analytical covariance matrix and with the “Template" initial parameters in Table 3 to generate the matrix and the MOPED compression, the biases in cosmological parameters are around $0.3\sigma$ ($0.1\sigma$) for the mean of the posterior (best-fit). Furthermore, the cosmological constraints with the analytical covariance matrix from the first iteration are also slightly tighter than the ones with the simulated covariance matrix.
@@
In either case, we then perform a second iteration, generating the analytical covariance matrix and the MOPED compression with the best-fit values from the first iteration. After the second iteration, table 4 shows the biggest deviation of $0.2\sigma$ arises in the constraint on the Hubble parameter with the analytical covariance matrix. Furthermore, if we also apply the MOPED compression, this is reduced to $0.1\sigma$. This applies regardless of whether we started with the better “Template” parameters or the “Guess” parameters. The mean of posterior for $\Omega_{b}h^{2}$ is generally shifted by around $0.14-0.18\sigma$. However, it is worth noting that the constraint on the baryonic density is quite tight because it is dominated by the BBN prior. We applied the BBN prior here because previous analysis shows spectroscopic clustering data has a very weak constraint on $\Omega_{b}h^{2}$. Since the shift in $\Omega_{b}h^{2}$ is much smaller than the width of the BBN prior, We believe this to be a sufficient demonstration of the method.
@@ 6 Conclusions
In this work, we combine the analytical covariance matrix, the MOPED compression, and the Taylor expansion interpolation of the power spectrum for the first time to analyze power spectrum data from the 6dFGS, BOSS, and eBOSS surveys. We demonstrate that this removes the necessity to generate large ensembles of simulations when analyzing such data (a potential bottleneck for future surveys such as the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2019) and Euclid (Laureijs et al., 2011)). while also offering potentially significant speed-ups in model fitting.
For the state-of-the-art dataset tested herein and without assuming a priori knowledge of the best-fit parameters, we use two different sets of initial parameters to generate the analytical covariance matrix and the MOPED compression. The “Template" parameters (within $1\sigma$ of the mean of the posterior with the simulated covariance matrix) and the “Guess" parameters (more than $3\sigma$ from the mean of the posterior with the simulated covariance matrix). The analytical covariance matrix takes around a day to compute, which is much faster than a few months for the simulated covariance matrix. Furthermore, the main bottleneck of the covariance matrix calculation is calculating the window kernels and their normalization in Wadekar & Scoccimarro (2020) from the random file, which is independent of the cosmology. The cosmology-dependent part of the calculation that needs to be recalculated for the second iteration only takes around half an hour. After two iterations, we found the mean of the posteriors and best-fit values of cosmological parameters from these two different sets agree with each other and are consistent with the ones with the simulated covariance matrix within $0.1-0.2\sigma$ regardless of the choice of initial guess. The remaining difference is likely due to the sampling noise.
We noticed the mean of the posterior of cosmological parameters does not necessarily agree with their best-fit values because of the prior volume effect. The best-fit values should be used to generate the analytical covariance matrix and the MOPED compression for the second iteration. Furthermore, the MOPED compression is able to speed up the MCMC by around a factor of four. This is because the matrix multiplication becomes the main bottleneck with the Taylor expansion. The MOPED compression can significantly reduce the size of the matrices which speeds up the MCMC.
Further work could investigate whether the compressed data vector itself could be emulated, further decreasing the run time. One could also determine a more rigorous set of convergence criteria for determining how many global iterations to perform, beyond the two we use here, that are suitable for guaranteeing the robustness of results from next-generation surveys